OpenStudy (anonymous):

substitution -x/6+y/2=1/2 x/3-y/6=-3/4

OpenStudy (anonymous):

help @waterineyes

OpenStudy (anonymous):

my y=-2x-9/2

OpenStudy (anonymous):

Multiply first by 2 and second by 12 and show me what you get..

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i.) -1/3X+Y=1 II.) 2X-2Y=-9

OpenStudy (anonymous):

Sorry..

OpenStudy (anonymous):

SORRY FOR?

OpenStudy (anonymous):

multiply first with 6 and second with 12 Finally you will get: -x + 3y = 3 4x - 2y = -9 Can you solve it now??

OpenStudy (anonymous):

thanks I will try

OpenStudy (anonymous):

x=3y-3

OpenStudy (anonymous):

If you now multiply 4 with first equation then by adding x will get cancel and you can easily find y.. Try on your own and let me know the values of x and y you calculated.. I am waiting..

OpenStudy (anonymous):

Yes: x = 3y - 3 Put it in second equation..

OpenStudy (anonymous):

Multiply first by 2 and second by 12 and show me what you get..

OpenStudy (anonymous):

then I will substitute it right?

OpenStudy (anonymous):

Yes this method is called \(\color{green}{SUBSTITUTION}\)..

OpenStudy (anonymous):

4x - 2y = -9 = 4(3y-3)-2y=-9 12y-2y=-9 10y=-9 y=-9/10?

OpenStudy (anonymous):

No check it once more..

OpenStudy (anonymous):

Where did -12 gone??

OpenStudy (anonymous):

12y-12-2y=-9

OpenStudy (anonymous):

Yes go ahead.

OpenStudy (anonymous):

12y-2y=-9+12

OpenStudy (anonymous):

Yes have faith just do the steps..

OpenStudy (anonymous):

10y=3 y=3/10?

OpenStudy (anonymous):

Yes now find x: x = 3y - 3

OpenStudy (anonymous):

x=3(3/10)-3 x=9/10-3 x=-21/10

OpenStudy (anonymous):

Yes, you have done it very correctly.. A medal for you from my side..

OpenStudy (anonymous):

Thanks a lot for your patients dear

OpenStudy (anonymous):

Its 1:26 am here.. May I go to sleep now?? My eyes are closing now..

OpenStudy (anonymous):

sure dear. Thanks for your help

OpenStudy (anonymous):

Welcome dear... Good Night nia.. Take care as well.. If you have more problems then I will clear them when I will wake up.. Try to solve more problems like this one you will surely get command over them..

OpenStudy (anonymous):

thanks a lot Dear

OpenStudy (anonymous):

Welcome..